# From Persona to Personalization: A Survey on Role-Playing Language Agents

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly boosted the rise of Role-Playing Language Agents (RPLAs), i.e., specialized AI systems designed to simulate assigned personas. By harnessing multiple advanced abilities of LLMs, including in-context learning, instruction following, and social intelligence, RPLAs achieve a remarkable sense of human likeness and vivid role-playing performance. RPLAs can mimic a wide range of personas, ranging from historical figures and fictional characters to real-life individuals. Consequently, they have catalyzed numerous AI applications, such as emotional companions, interactive video games, personalized assistants and copilots, and digital clones. In this paper, we conduct a comprehensive survey of this field, illustrating the evolution and recent progress in RPLAs integrating with cutting-edge LLM technologies. We categorize personas into three types: 1) Demographic Persona, which leverages statistical stereotypes; 2) Character Persona, focused on well-established figures; and 3) Individualized Persona, customized through ongoing user interactions for personalized services. We begin by presenting a comprehensive overview of current methodologies for RPLAs, followed by the details for each persona type, covering corresponding data sourcing, agent construction, and evaluation. Afterward, we discuss the fundamental risks, existing limitations, and future prospects of RPLAs. Additionally, we provide a brief review of RPLAs in AI applications, which reflects practical user demands that shape and drive RPLA research. Through this work, we aim to establish a clear taxonomy of RPLA research and applications, and facilitate future research in this critical and ever-evolving field, and pave the way for a future where humans and RPLAs coexist in harmony.
@@ 1 Introduction
Digital life has been a pursuit for humanity for decades, reflecting our deep-rooted fascination with the intersection of technology and human experience. Bridging this pursuit with imaginative concepts, role-playing AI systems embody the digital life by bringing these personas to life in interactive forms. These systems, which simulate assigned personas, have long been a concept in the human imagination, capturing the essence of our desire to create and interact with artificial beings that can understand, respond, and engage with us in a seemingly sentient manner. With role-playing agents, various personas can be replicated by their agent counterparts, including historical figures, fictional characters, or individuals in our daily lives. Recently, focusing on the text modality, Role-Playing Language Agents (RPLAs) are coming into reality (Shanahan et al., 2023; Shao et al., 2023; Wang et al., 2024c), which inspires a wide range of novel applications, such as digital clones for individuals (Ng et al., 2024), AI characters in chatbots (Wang et al., 2023g), and role-playing video games (Wang et al., 2023a), even stimulating social science research (Rao et al., 2023). As RPLAs become increasingly integrated into our daily lives, it is essential to foster a society that thrives on the synergistic coexistence of humans and these intelligent agents.
Recent developments in Large Language Models (LLMs)  (OpenAI, 2023; Google, 2023; Anthropic, 2024) have greatly facilitated the emergence of RPLAs. LLMs grow adept at producing a compelling sense of human likeness (Shanahan et al., 2023; Zhou et al., 2024b), and can be regarded as superpositions of beliefs (Kovač et al., 2023) and personas (Lu et al., 2024). Furthermore, with alignment training, LLMs are able to adhere to the instruction of persona role-playing, including replicating their knowledge (Lu et al., 2024; Li et al., 2023a), linguistic and behavior patterns (Wang et al., 2023g; Zhou et al., 2023a), and even underlying personalities (Shao et al., 2023; Wang et al., 2024c). They are able to both mimic the personas as prompted in the contexts (Wang et al., 2023g; Li et al., 2023a), or harness their inherent parametric knowledge for widely-recognized demographics or characters (Shao et al., 2023; Lu et al., 2024). Considering their practical significance, there has been an increase in research efforts dedicated to RPLAs with LLMs, including their development (Wang et al., 2023g; Li et al., 2023a; Zhou et al., 2023a), analysis (Shao et al., 2023; Yuan et al., 2024b), and applications (Rao et al., 2023; Park et al., 2023; Mysore et al., 2023). Conversely, RPLAs also offer an ideal perspective and testing ground for investigating the behaviors and capabilities of LLMs and language agents, particularly those related to social interactions (Li et al., 2023d; Chen et al., 2023a).
@@
    Character Persona, which represents well-established and widely-recognized individuals, especially in the existing literature, including celebrities, historical figures, and fictional characters. Role-playing these personas challenges models’ capability in understanding curated materials of the existing characters, harnessing knowledge in LLMs’ parameters or given contexts (Shao et al., 2023; Wang et al., 2023g; 2024c).
@@ 2.1 The Roadmap of Large Language Models
Several key abilities have emerged in LLMs (Wei et al., 2022a) throughout their evolution, including in-context learning (Brown et al., 2020), instruction following (Ouyang et al., 2022), step-by-step reasoning (Wei et al., 2022b), and social intelligence (Wang et al., 2024a; Sclar et al., 2023; Light et al., 2023), which lay the foundation for complicated role-playing behavior of LLMs towards RPLAs. First, the in-context learning ability allows LLMs to learn information from prompts without parameter updates. This facilitates LLMs’ adaptation to the provided knowledge of various characters and mimicking their behaviors by following example demonstrations. Second, the instruction following ability enables LLMs to adhere to role-playing instructions, such as “Serve as a helpful assistant” or “Role-play Hermione Granger in the Harry Potter Series. <Description>. <Example Conversations>. <Requirements>.”. Finally, step-by-step reasoning and social intelligence refine LLMs in terms of anthropomorphic cognition, contributing to an enriched sense of human likeness and nuanced emotional support in RPLA applications.
@@ 2.2 LLM-powered Language Agents
The AI community has long been pursuing the concept of “agent”, approaching the intelligence and autonomy of humans. Traditional symbolic agents (Bernstein, 2001; Küngas et al., 2004) and reinforce-learning agents (Fachantidis et al., 2017; Florensa et al., 2018) mainly optimize their actions based on rules or pre-defined rewards. Research in language agents primarily focuses on training within constrained environments with limited knowledge, diverging from the complex and diverse nature of the human learning process. However, such agents struggle to emulate complicated human-like behaviors, particularly in open-domain settings (Mnih et al., 2015; Lillicrap et al., 2015; Schulman et al., 2017; Haarnoja et al., 2017). Recently, LLMs have demonstrated remarkable capabilities with promising potential in achieving human-level intelligence, which has sparked a rise in research focusing on LLM-based language agents (Sclar et al., 2023; Chalamalasetti et al., 2023; Liu et al., 2023d; Xie et al., 2024a). Research in this area primarily involves equipping LLMs with essential human-like capabilities, such as planning, tool-usage and memory (Weng, 2023), which are essential for developing advanced RPLAs with anthropomorphic cognition and abilities.
@@
In many real-world scenarios, the agents need to make long-horizon planning to solve complex tasks (Rana et al., 2023; Yuan et al., 2023). When facing these tasks, LLM-powered agents could decompose the complex tasks into subtasks and adopt various planning strategies, e.g., CoT (Wei et al., 2022b) and ReAct (Yao et al., 2023b), to adaptively plan for the next action with feedback from environments (Wang et al., 2023a; Gotts et al., 2003; Wang et al., 2023j; Song et al., 2023; Zhang et al., 2024b). For RPLAs, these adaptive planning strategies enable them to simulate realistic and dynamic interactions in complicated environments such as games (Wang et al., 2023a) and social simulations (Park et al., 2023).
@@
The memory mechanism stores the profile of agents along with environmental information to assist agents in future actions. The profile typically includes basic information (age, gender, career), psychological traits (reflecting personality), and social relationships (Wang et al., 2023c; Park et al., 2023; Qian et al., 2023), which can be manually created (Caron & Srivastava, 2022; Zhang et al., 2023a; Pan & Zeng, 2023; Huang et al., 2023b; Karra et al., 2022; Safdari et al., 2023) or generated from models (Wang et al., 2023c). This module enables agents to accumulate experiences, evolve, and act consistently and effectively (Park et al., 2023). Language agents draw from cognitive science research on human memory, which progresses from sensory to short-term, then to long-term memory (Atkinson & Shiffrin, 1968; Craik & Jennings, 1992). The short-term memory is regarded as the information input within the constraint window of transformer architecture (Fischer, 2023; Rana et al., 2023; Wang et al., 2023j; Zhu et al., 2023a). In contrast, long-term memory is usually reserved in the external vector storage (Qian et al., 2023; Zhong et al., 2023; Zhu et al., 2023b; Lin et al., 2023; Xie et al., 2023; Wu et al., 2024b) or natural languages database (Shinn et al., 2023; Modarressi et al., 2023), from which agents can quickly query and retrieve information as required. Compared to vanilla LLMs, language agents need to learn and perform tasks in changing environments. For RPLAs, the memory mechanism plays a pivotal role by enabling these agents to maintain continuity and context in interactions over time. By storing and retrieving user-specific data and environmental context, agents deliver more personalized and relevant responses, thus enhancing user experience and engagement in diverse scenarios.
@@
for tree= forked edges, grow’=0, calign=child edge, calign child=(n_children()+1)/2, , [RPLAs, root [Demographic Persona §4, demographic [Analysis, demographic [Inherent Demographics, demographic [Serapio-García et al. (2023); MPI; PsychoBench; OpinionQA; Hartmann et al. (2023), demographic_work] ] [Demographic Role-playing, demographic [Chain Prompting; Serapio-García et al. (2023); Deshpande et al. (2023); PAIR, demographic_work] ] ] [Application, demographic [Task Sovling, demographic [ MetaGPT; OKR-AGENT, demographic_work ] ] [Social Behaviors, demographic [ Cicero; The Avalon; The Werewolf ; Humaniod Agent , demographic_work ] ] ] ] [Character Persona §5, character [Data Source, character [ChatHaruhi; CharacterGLM; Character-LLM; RoleLLM; HPD; CharacterEval; DITTO; NarrativePlay; MORTISE, character_work] ] [Construction, character [Parametric Training, character [Character-LLM; ChatHaruhi; Neeko; RoleLLM, character_work] ] [Nonparametric Prompting, character [RoleLLM; CharacterGLM; ChatHaruhi; InCharacter; LifeChoice, character_work] ] ] [Evaluation, character [Character-independent Capabilities, character [ Character-LLM; InCharacter; CharacterGLM; CharacterEval; Deshpande et al. (2023); Kosinski (2023); Sorin et al. (2023); Huang et al. (2023a); Zhou et al. (2024b); Wang et al. (2024a), character_work ] ] [ Linguistic Style and Knowledge, character [RoleLLM; ChatHaruhi; CharacterGLM; Neeko; HPD; Character-LLM; RoleEval; NarrativePlay; MORTISE; DITTO, character_work] ] [ Personality and Thinking Process, character [Character-LLM; LifeChoice; RoleInteract; InCharacter; Cross-MR, character_work] ] ] ] [Individualized Persona §6, personalization [Data Source, personalization [Crowd-sourcing, personalization [Zhang et al. (2018); ConvAI2; FoCus; RealPersonaChat; KBP; Cho et al. (2023); ECHO, personalization_work] ] [Internet, personalization [P-Ubuntu; MPCHAT; COMSET; LiveChat, personalization_work] ] [LLM, personalization [Synthetic-Persona-Chat, personalization_work] ] ] [Construction, personalization [Offline Learning, personalization [PersonaAware-D2S; GIRL; personalized-RLHF;  Jang et al. (2023); VaRMI; ControlLM, personalization_work] ] [Online Learning, personalization [TeachMe;  Kim et al. (2024);  Baek et al. (2023); CoPS; PEARL; Persona-DB; ExploreLLM;  Qin et al. (2024); GeneInput;  Bai et al. (2022);ECHO, personalization_work] ] ] [Evaluation, personalization [Conversation, personalization [LiveChat; FoCus; Poly-encoders; MPCHAT; PRM;  Dinan et al. (2018);ECHO , personalization_work] ] [Recommend-ation, personalization [LFM; PALR; IGL; ONCE; PEARL;  Dai et al. (2023a);  Hou et al. (2024);  Kang et al. (2023);  Liu et al. (2023a), personalization_work] ] [Task Solving, personalization [G4C; MMToM-QA; OpenToM; Toolllm; Api-bank; Toolqa; UltraTool; AutoDroid; TaskBench; Assistgui; Planbench , personalization_work]] ] ] [Risks §7, risk [Toxicity, risk [PDP;  Deshpande et al. (2023); MPCHAT,risk_work] ] [Bias, risk [PROMPTINJECT; Branch et al. (2022);  Zheng et al. (2023a); Kotek et al. (2023); Marked Personas; CAMEL; Rutinowski et al. (2023); Hartmann et al. (2023); Xue et al. (2023); Coda-Forno et al. (2023), risk_work] ] [Hallucination, risk [ Li et al. (2016); Zhang et al. (2018); Neel et al. (2021); Pawelczyk et al. (2023); SPP; Character-LLM, risk_work] ] [Privacy Violations, risk [K-Anonymity; ProPILE;  Li et al. (2023e); Staab et al. (2023), risk_work] ] [Technical Challenge in Real-world Deployment, risk[ COSPLAY; FANToM; Character-LLM;  Liu et al. (2023b); Shao et al. (2023); CharacterEval, risk_work] ] ] ]

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have significantly boosted the rise of Role-Playing Language Agents (RPLAs), i.e., specialized AI systems designed to simulate assigned personas. By harnessing multiple advanced abilities of LLMs, including in-context learning, instruction following, and social intelligence, RPLAs achieve a remarkable sense of human likeness and vivid role-playing performance. RPLAs can mimic a wide range of personas, ranging from historical figures and fictional characters to real-life individuals. Consequently, they have catalyzed numerous AI applications, such as emotional companions, interactive video games, personalized assistants and copilots, and digital clones. In this paper, we conduct a comprehensive survey of this field, illustrating the evolution and recent progress in RPLAs integrating with cutting-edge LLM technologies. We categorize personas into three types: 1) Demographic Persona, which leverages statistical stereotypes; 2) Character Persona, focused on well-established figures; and 3) Individualized Persona, customized through ongoing user interactions for personalized services. We begin by presenting a comprehensive overview of current methodologies for RPLAs, followed by the details for each persona type, covering corresponding data sourcing, agent construction, and evaluation. Afterward, we discuss the fundamental risks, existing limitations, and prospects of RPLAs. Additionally, we provide a brief review of RPLAs in AI products in the market, which reflects practical user demands that shape and drive RPLA research. Through this survey, we aim to establish a clear taxonomy of RPLA research and applications, facilitate future research in this critical and ever-evolving field, and pave the way for a future where humans and RPLAs coexist in harmony.
@@ 1 Introduction
Digital life has been a pursuit for humanity for decades, reflecting our deep-rooted fascination with the intersection of technology and human experience. Bridging this pursuit with imaginative concepts, role-playing AI systems embody the digital life by bringing these personas to life in interactive forms. These systems, which simulate assigned personas, have long been a concept in the human imagination, capturing the essence of our desire to create and interact with artificial beings that can understand, respond, and engage with us in a seemingly sentient manner. With role-playing agents, various personas can be replicated by their agent counterparts, including historical figures, fictional characters, or individuals in our daily lives. Recently, focusing on the text modality, Role-Playing Language Agents (RPLAs) are coming into reality (Shanahan et al., 2023; Shao et al., 2023; Wang et al., 2024d), which inspires a wide range of novel applications, such as digital clones for individuals (Xu et al., 2024b; Ng et al., 2024), AI characters in chatbots (Wang et al., 2024a), and role-playing video games (Wang et al., 2023a), even stimulating social science research (Rao et al., 2023). As RPLAs become increasingly integrated into our daily lives, it is essential to foster a society that thrives on the synergistic coexistence of humans and these intelligent agents.
Recent developments in Large Language Models (LLMs)  (OpenAI, 2023; Google, 2023; Anthropic, 2024) have greatly facilitated the emergence of RPLAs. LLMs grow adept at producing a compelling sense of human likeness (Shanahan et al., 2023; Zhou et al., 2024b), and can be regarded as superpositions of beliefs (Kovač et al., 2023) and personas (Lu et al., 2024). Furthermore, with alignment training, LLMs are able to adhere to the instruction of persona role-playing, including replicating their knowledge (Lu et al., 2024; Li et al., 2023a), linguistic and behavior patterns (Wang et al., 2024a; Zhou et al., 2023a), and even underlying personalities (Shao et al., 2023; Wang et al., 2024d). They are able to both mimic the personas as prompted in the contexts (Wang et al., 2024a; Li et al., 2023a), or harness their inherent parametric knowledge for widely-recognized demographics or characters (Shao et al., 2023; Lu et al., 2024). Considering their practical significance, there has been an increase in research efforts dedicated to RPLAs with LLMs, including their development (Wang et al., 2024a; Li et al., 2023a; Zhou et al., 2023a), analysis (Shao et al., 2023; Yuan et al., 2024b), and applications (Rao et al., 2023; Park et al., 2023; Mysore et al., 2023). Conversely, RPLAs also benefit the development of LLMs and language agents. They offer an ideal perspective and testing ground for investigating the behaviors and capabilities of LLMs and language agents, particularly those related to social interactions (Li et al., 2023d; Chen et al., 2023a; Wu et al., 2024c). They also facilitate the creation of diverse and massive synthetic data for LLM training at scale (Chan et al., 2024).
@@
    Character Persona, which represents well-established and widely-recognized individuals, especially in the existing literature, including celebrities, historical figures, and fictional characters. Role-playing these personas challenges models’ capability in understanding curated materials of the existing characters, harnessing knowledge in LLMs’ parameters or given contexts (Shao et al., 2023; Wang et al., 2024a; d).
@@ 2.1 The Roadmap of Large Language Models
Several key abilities have emerged in LLMs (Wei et al., 2022a) throughout their evolution, including in-context learning (Brown et al., 2020), instruction following (Ouyang et al., 2022), step-by-step reasoning (Wei et al., 2022b), and social intelligence (Wang et al., 2024b; Sclar et al., 2023; Light et al., 2023), which lay the foundation for complicated role-playing behavior of LLMs towards RPLAs. First, the in-context learning ability allows LLMs to learn information from prompts without parameter updates. This facilitates LLMs’ adaptation to the provided knowledge of various characters and mimicking their behaviors by following example demonstrations. Second, the instruction following ability enables LLMs to adhere to role-playing instructions, such as “Serve as a helpful assistant” or “Role-play Hermione Granger in the Harry Potter Series. <Description>. <Example Conversations>. <Requirements>.”. Finally, step-by-step reasoning and social intelligence refine LLMs in terms of anthropomorphic cognition, contributing to an enriched sense of human likeness and nuanced emotional support in RPLA applications.
@@ 2.2 LLM-powered Language Agents
The AI community has long been pursuing the concept of “agent”, approaching the intelligence and autonomy of humans. Traditional symbolic agents (Bernstein, 2001; Küngas et al., 2004) and reinforce-learning agents (Fachantidis et al., 2017; Florensa et al., 2018) mainly optimize their actions based on rules or pre-defined rewards. Research in language agents primarily focuses on training within constrained environments with limited knowledge, diverging from the complex and diverse nature of the human learning process. However, such agents struggle to emulate complicated human-like behaviors, particularly in open-domain settings (Mnih et al., 2015; Lillicrap et al., 2015; Schulman et al., 2017; Haarnoja et al., 2017). Recently, LLMs have demonstrated remarkable capabilities with promising potential in achieving human-level intelligence, which has sparked a rise in research focusing on LLM-based language agents (Sclar et al., 2023; Chalamalasetti et al., 2023; Liu et al., 2023d; Xie et al., 2024b). Research in this area primarily involves equipping LLMs with essential human-like capabilities, such as planning, tool-usage and memory (Weng, 2023), which are essential for developing advanced RPLAs with anthropomorphic cognition and abilities.
@@
In many real-world scenarios, the agents need to make long-horizon planning to solve complex tasks (Rana et al., 2023; Yuan et al., 2023). When facing these tasks, LLM-powered agents could decompose the complex tasks into subtasks and adopt various planning strategies, e.g., CoT (Wei et al., 2022b) and ReAct (Yao et al., 2023b), to adaptively plan for the next action with feedback from environments (Wang et al., 2023a; Gotts et al., 2003; Wang et al., 2023i; Song et al., 2023; Zhang et al., 2024b). For RPLAs, these adaptive planning strategies enable them to simulate realistic and dynamic interactions in complicated environments such as games (Wang et al., 2023a) and social simulations (Park et al., 2023).
@@
The memory mechanism stores the profile of agents along with environmental information to assist agents in future actions. The profile typically includes basic information (age, gender, career), psychological traits (reflecting personality), and social relationships (Wang et al., 2023c; Park et al., 2023; Qian et al., 2023), which can be manually created (Caron & Srivastava, 2022; Zhang et al., 2023a; Pan & Zeng, 2023; Huang et al., 2023b; Karra et al., 2022; Safdari et al., 2023) or generated from models (Wang et al., 2023c). This module enables agents to accumulate experiences, evolve, and act consistently and effectively (Park et al., 2023). Language agents draw from cognitive science research on human memory, which progresses from sensory to short-term, then to long-term memory (Atkinson & Shiffrin, 1968; Craik & Jennings, 1992). The short-term memory is regarded as the information input within the constraint window of transformer architecture (Fischer, 2023; Rana et al., 2023; Wang et al., 2023i; Zhu et al., 2023a). In contrast, long-term memory is usually reserved in the external vector storage (Qian et al., 2023; Zhong et al., 2023; Zhu et al., 2023b; Lin et al., 2023; Xie et al., 2023; Wu et al., 2024b) or natural languages database (Shinn et al., 2023; Modarressi et al., 2023), from which agents can quickly query and retrieve information as required. Compared to vanilla LLMs, language agents need to learn and perform tasks in changing environments. For RPLAs, the memory mechanism plays a pivotal role by enabling these agents to maintain continuity and context in interactions over time. By storing and retrieving user-specific data and environmental context, agents deliver more personalized and relevant responses, thus enhancing user experience and engagement in diverse scenarios.
@@
for tree= forked edges, grow’=0, calign=child edge, calign child=(n_children()+1)/2, , [RPLAs, root [Demographic Persona §4, demographic [Analysis, demographic [Inherent Demographics, demographic [Serapio-García et al. (2023); MPI; PsychoBench; OpinionQA; Hartmann et al. (2023), demographic_work] ] [Demographic Role-playing, demographic [Chain Prompting; Serapio-García et al. (2023); Deshpande et al. (2023); PAIR, demographic_work] ] ] [Application, demographic [Task Solving, demographic [ MetaGPT; OKR-AGENT, demographic_work ] ] [Social Behaviors, demographic [ Cicero; The Avalon; The Werewolf ; Humaniod Agent , demographic_work ] ] ] ] [Character Persona §5, character [Data Source, character [ChatHaruhi; CharacterGLM; Character-LLM; RoleLLM; HPD; CharacterEval; DITTO; NarrativePlay; MORTISE; MMRole;TimeChara;PersonaHub, character_work] ] [Construction, character [Parametric Training, character [Character-LLM; ChatHaruhi; Neeko; RoleLLM, character_work] ] [Nonparametric Prompting, character [RoleLLM; CharacterGLM; ChatHaruhi; InCharacter; LifeChoice, character_work] ] ] [Evaluation, character [Character-independent Capabilities, character [ Character-LLM; InCharacter; CharacterGLM; CharacterEval; MMRole;PersonaGym;TimeChara; Deshpande et al. (2023); Kosinski (2023); Sorin et al. (2023); Huang et al. (2023a); Zhou et al. (2024b); Wang et al. (2024b), character_work ] ] [ Linguistic Style and Knowledge, character [RoleLLM; ChatHaruhi; CharacterGLM; Neeko; HPD; Character-LLM; RoleEval; NarrativePlay; MORTISE; DITTO, character_work] ] [ Personality and Thinking Process, character [Character-LLM; LifeChoice; SocialBench; InCharacter; Cross-MR, character_work] ] ] ] [Individualized Persona §6, personalization [Data Source, personalization [Crowd-sourcing, personalization [Zhang et al. (2018); ConvAI2; FoCus; RealPersonaChat; KBP; Cho et al. (2023); ECHO, personalization_work] ] [Internet, personalization [P-Ubuntu; MPCHAT; COMSET; LiveChat, personalization_work] ] [LLM, personalization [Synthetic-Persona-Chat, personalization_work] ] ] [Construction, personalization [Offline Learning, personalization [PersonaAware-D2S; GIRL; personalized-RLHF;  Jang et al. (2023); VaRMI; ControlLM, personalization_work] ] [Online Learning, personalization [TeachMe;  Kim et al. (2024);  Baek et al. (2023); CoPS; PEARL; Persona-DB; ExploreLLM;  Qin et al. (2024); GeneInput;  Bai et al. (2022a);ECHO, personalization_work] ] ] [Evaluation, personalization [Conversation, personalization [LiveChat; FoCus; Poly-encoders; MPCHAT; PRM;  Dinan et al. (2018);ECHO , personalization_work] ] [Recommend-ation, personalization [LFM; PALR; IGL; ONCE; PEARL;  Dai et al. (2023a);  Hou et al. (2024);  Kang et al. (2023);  Liu et al. (2023a), personalization_work] ] [Task Solving, personalization [G4C; MMToM-QA; OpenToM; Toolllm; Api-bank; Toolqa; UltraTool; AutoDroid; TaskBench; Assistgui; Planbench , personalization_work]] ] ] [Risks §7, risk [Toxicity, risk [PDP;  Deshpande et al. (2023); MPCHAT,risk_work] ] [Bias, risk [PROMPTINJECT; Branch et al. (2022);  Zheng et al. (2023a); Kotek et al. (2023); Marked Personas; CAMEL; Rutinowski et al. (2023); Hartmann et al. (2023); Xue et al. (2023); Coda-Forno et al. (2023), risk_work] ] [Hallucination, risk [ Li et al. (2016); Zhang et al. (2018); Neel et al. (2021); Pawelczyk et al. (2023); SPP; Character-LLM, risk_work] ] [Privacy Violations, risk [K-Anonymity; ProPILE;  Li et al. (2023e); Staab et al. (2023), risk_work] ] [Technical Challenge in Real-world Deployment, risk[ COSPLAY; FANToM; Character-LLM;  Liu et al. (2023b); Shao et al. (2023); CharacterEval, risk_work] ] ] ]